### v3

# Kähler Solitons, Contact Structures, and Isoparametric Functions

Abstract.
All known examples of gradient Kähler-Ricci soliton in real dimension four are toric and the symmetry is intrinsically related to the potential function $f$ and the scalar curvature $S\SS$. In this article, we consider the case that $f$ and $S\SS$ are functionally dependent and deduce a complete classification, while the independence case is addressed elsewhere. The main theorem recovers all known examples of cohomogeneity one symmetry. We also discover a connection to the theory of isoparametric functions and contact geometry. Indeed, a key ingredient is a new characterization for a deformed Sasakian structure generalizing a classical result.
Department of Mathematics and Statistics, Texas Tech University, Lubbock, TX 79409, Email: hung.tran@ttu.edu
1. Introduction
A gradient Ricci soliton (GRS) $(M,g,f,\lambda)$ is a Riemannian manifold with metric $g$, potential function $f$, and a constant $\lambda$ such that, for $\mathrm{Rc}$} denoting the Ricci curvature,
(1.1) ${\mathrm{Rc}}+\text{Hess}{f}=\lambda g.$
Such a structure is a self-similar solution to the Ricci flow and plays a crucial role in its analysis. The general theory was introduced by R. Hamilton [40] and has several celebrated applications including [53, 54, 11, 14, 15, 20]. Recently, there have been significant breakthroughs towards a Ricci flow program in higher dimensions [13, 4, 3, 2, 66, 63, 33, 41, 39]. Dimension four seems to be the most tractable and it is of tremendous interest to investigate corresponding GRS.
A gradient Kähler Ricci soliton (GKRS) $(M,g,f,J,\lambda)$ is a GRS such that $(M,g,J)$ is Kähler for a complex structure $J$. In particular, the soliton vector field $\nabla f$ is automatically holomorphic. The subject has an extensive literature and recent efforts lead to the classification of all GKRS surfaces with $\lambda>0$ [29, 28, 5, 46]. Thus, there is realistic hope to obtain a full classification of all GKRS in real dimension four. Indeed, all known examples have toric symmetry which is intrinsically related to the potential function $f$ and the scalar curvature $S\SS$. Here we propose a strategy based on the relation between these functions and applicable to all signs of $\lambda$. It turns out that there will be a dichotomy as described below.
Definition 1.1.
Two functions are functionally dependent if their gradients are scalar multiples of each other at points they are non-zero. Otherwise, they are functionally independent.
In a companion article [69], we consider the independence case and prove that the soliton admits a torus action under a generic assumption. In this paper, we focus on the dependence case and deduce a complete classification. Indeed, on a GRS, $f$ and $S\SS$ are functionally dependent if and only if $\Delta f$ and $|\nabla f|$ are constant on each connected component of a level set of $f$ (see Proposition 2.11). This property generally defines an isoparametric function; the study of such functions in space forms was motivated by questions in geometric optics [62, 21, 45]. The classification in an ambient round sphere, Question 34 in S. T. Yau’s list [71], is remarkably deep and has attracted enormous interest by, for example, [50, 49, 1, 23, 26]. Certain aspects of the theory is extended to a general Riemannian manifold [70, 38, 48, 57, 31] and will contribute to our analysis.
In the study of a GKRS, the early constructions of examples with maximal symmetry [44, 17, 24, 18, 35, 52], all have isoparametric potential functions. Without the Kähler setup, an equivalent condition is considered in [42]. It also arises as a consequence of constant scalar curvature [36, 25]. Furthermore, in [29, Section 3], the authors discovered that an appropriate rescaling of the potential function, on an expanding GKRS, converges to an isoparametric function with respect to a suitable metric. Our first result classifies complete Kähler GRS surfaces with such a potential function.
Theorem A.
Let $(M,g,J,f)$ be a complete connected non-flat Kähler GRS of real dimension four. If $f$ and $S\SS$ are functionally dependent then the manifold must be either
  • a product of a constant curvature surface with a $2D$ Kähler GRS, or
  • is foliated by equidistant level sets of $f$ and the soliton equation (1.1) reduces to ODEs which can be solved explicitly. Each regular level set is a quotient, by a discrete group, of a deformed Sasakian space-form in dimension three and there is at least one singular level set.
A soliton structure on a manifold can always be lifted to its universal cover and we have a refined statement.
Corollary B.
Let $(M,g,J,f)$ be a simply connected irreducible non-compact GKRS of real dimension four. If $f$ and $S\SS$ are functionally dependent then the manifold must be of cohomogeneity one. Each principal orbit is a connected homogeneous Sasakian structure with constant holomorphic sectional curvature. The singular set corresponds to a singular orbit of the cohomogeneity one action and is simply-connected. Consequently, the GKRS is of maximal irreducible symmetry (isometry group of dimension four).
Remark 1.2.
See [68] for the classification of GKRS with such symmetry. In particular, the cohomogeneity one action is given by the automorphism group of the Sasakian structure. There are three models of simply connected Sasakian structures with constant holomorphic sectional curvature [65]. Furthermore, if $\lambda\geq 0$, then the Sasakian structure must be on a sphere or its quotient along the Hopf fibration by a cyclic group, the isometry group is $U(2)$, and the singular orbit is a point or round $\mathbb{CP}^{1}$. Then, $M$ is either $\mathbb{C}^{2}$ or a holomorphic line bundle $L^{\ell}$, $\ell\in\mathbb{Z}$, over $\mathbb{CP}^{1}$. $\ell$ corresponds to each principal orbit being the quotient by $\mathbb{Z}_{|\ell|}$ of the spherical Sasakian structure and also denotes the first Chern class of the bundle. By [35, Lemma 1.2], one obtains a shrinking/steady/expanding soliton depending on whether $\ell$ is greater/equal/less than 2-2. If $\lambda<0$, then each Sasakian space-form could arise and the analysis is given in [30, Theorem 4.20]. Hence, the classification includes all $U(2)$-invariant metrics constructed by [44, 17, 24, 18, 35, 52].
In relation to the recent classification of shrinking GKRS in real dimension four, our assumption recovers all models except for $T^{2}$-symmetry metrics on $Bl_{2}(\mathbb{CP}^{2})$ and $Bl_{1}(\mathbb{C}^{1}\times\mathbb{CP}^{1})$. They correspond to the case $f$ and $S\SS$ are functionally independent.
Remark 1.3.
The dimension assumption is crucial as [30] describes a more general construction in higher dimensions such that $f$ and $S\SS$ are functionally dependent.
Furthermore, there is a partial converse.
Corollary C.
Let $(M,g,J,f)$ be a complete connected irreducible Kähler GRS surfaces of cohomogeneity one in real dimension four. If either $\lambda\neq 0$ or the isometry group is compact, then $f$ and $S\SS$ are functionally dependent and each principal orbit is a connected deformed homogeneous Sasakian structure.
The results above suggest a connection between a GKRS and a Sasakian structure, a fundamental notion of contact geometry. This arguably is an odd-dimensional counterpart of symplectic geometry and both lie at the heart of classical mechanics. Thus, it has rich literature; for example, contact transformation was studied by S. Lie [47]. Here we focus on a direction with more attention to an associated Riemannian metric. The current theory owns much to the foundational work of S. Sasaki, D. Blair, and others [60, 59, 61, 8, 9]; see also a recent book [12] and a survey [10].
An odd-dimensional Riemannian manifold $(P,g_{P})$- with a vector field $\zeta$, a $1$-form $\eta$, and a tensor field of type $(1,1)$ $\Phi$- is called an almost contact metric structure if
$\eta(\zeta)=1,~~\Phi^{2}=-\mathrm{Id}+\zeta\otimes\eta,\text{ and }g_{P}(\Phi(X),\Phi(Y))=g_{P}(X,Y)-\eta(X)\eta(Y).$
An almost contact metric structure $(P,\zeta,\eta,\Phi,g)$ is called a contact metric structure if
$d\eta(X,Y)=g(X,\Phi(Y)).$
Moreover, a Sasakian structure is a contact metric such that the cone $(M\times\mathbb{R}^{+},r^{2}g+dr^{2})$ is Kähler. A deformed contact/Sasakian structure is obtained by a specific rescaling of the Reeb vector field and the transverse metric; see Definition 2.7.
Let $(M,g,J,f)$ be a Kähler GRS. Let $M_{c}$ be a regular level set of $f$ and
(1.2) $V:$ $=\frac{\nabla f}{|\nabla f|},~~g_{c}:=g_{\mid M_{c}},~~\zeta_{c}:=-J(V),~~\eta_{c}(\cdot):=g(\cdot,\zeta_{c}),~~\Phi_{c}(\cdot):=-\eta_{c}(\cdot)V+J(\cdot).$
Together, $(M_{c},\zeta_{c},\eta_{c},\Phi_{c},g_{c})$ is an almost contact metric structure. Our next theorem shows that going from almost contact to deformed contact imposes significant restriction on the soliton structure leading to a full classification.
Theorem D.
Let $(M,g,J,f,\lambda)$ be a complete connected Kähler GRS. For each regular value $c$, suppose that $(M_{c},\zeta_{c},\eta_{c},\Phi_{c},g_{c})$ is a deformed contact structure. Then the soliton is totally determined by a connected Sasakian model $(P,\eta,\zeta,\Phi,g_{P})$ which is a Riemannian submersion, with circle fibers, over a Kähler-Einstein manifold $(N,g_{N},J_{N})$ with ${\mathrm{Rc}}_{N}=kg_{N}$. That is, there is submersion map $\pi:P\mapsto N$ such that
$g_{P}=\eta\otimes\eta+\pi^{\ast}g_{N},~~~d\eta=\pi^{\ast}\omega_{N}.$
There is an interval $I$ with coordinate $s$ such that there is a diffeomorphism $\phi:I\times P\mapsto M_{o}$, a dense subset of $M$, $f\circ\phi=Bs+C$, and
(1.3) $\phi^{\ast}g=\frac{ds^{2}}{\alpha(s)}+\alpha(s)\eta\otimes\eta+(2s+A)\pi^{\ast}g_{N}.$
Here $A,B,C$ are constant and $\alpha$ solves a first order equation, for $n=\dim_{\mathbb{C}}N$, namely
$\lambda(2s+A)=k-\frac{d\alpha}{ds}-\frac{2n\alpha}{2s+A}+B\alpha.$
There is a boundary point of $I$ such that $\alpha\rightarrow 0$. Furthermore, if $(2s+A)\rightarrow 0$ towards that end point, then $(P,\eta,\zeta,\Phi,g_{P})$ is the standard Sasakian sphere and $(N,g_{N},J_{N})$ is, up to homothety, isomorphic to a standard complex projective space.
Remark 1.4.
The main content of the theorem is about $|\nabla f|$ constant on each connected component of a level set of $f$ (rectifiable), constant Ricci, and the explicit Sasakian structure. Equation (1.3) is a variant of the Calabi ansatz [16]. There are also other studies which construct Kähler GRS from Sasaki-Einstein manifolds [37, 10].
Indeed, the proofs of Theorem A and D both rely on understanding a deformed contact structure. Thus, the following result might be of independent interest.
Theorem E.
An almost contact metric manifold $(P,\zeta,\eta,\Phi,g)$ satisfies
$(\nabla_{X}\Phi)(Y)=bg(X,Y)\zeta-\eta(Y)bX.$
for every vector fields $X$ and $Y$ if and only if
  1. (i)
    for $b\neq 0$, it is a deformed Sasakian structure;
  2. (ii)
    for $b=0$, $\zeta$ is parallel and the universal cover is a product of a line or circle with a Kähler manifold whose almost complex structure is induced by $\Phi$.
Remark 1.5.
The case $b=1$ is a classical result [12, Theorem 7.3.16].
Theorem E has an important consequence. Let $(M,g,J)$ be a Kähler manifold and $P$ a real hypersurface with unit normal $V$ and shape operator $LX=\nabla_{X}V.$ Let $\zeta=-JV$, $\eta$ the dual 1-form, and $\Phi(\cdot):=J(\cdot)-\eta(\cdot)V.$ It is immediate that $(P,\zeta,\eta,\Phi,g)$ is an almost contact metric structure.
Theorem F.
$L=\alpha\text{Id}+\beta\zeta\otimes\eta$ for a constant $\alpha$ iff $(P,\eta,\zeta,\Phi,g)$ is
  1. (i)
    for $\alpha\neq 0$, a deformed-Sasakian structure.
  2. (ii)
    for $\alpha=0$, locally a Riemannian product.
The organization of the paper is as follows. Section 2 will recall definitions and preliminary results. The proofs of Theorems E and F will be given in Section 3: using the theory of foliation and toolkit developed by B. O’Neill [51] to track how the covariant derivative of $\Phi$ changes under a deformation. Section 4 investigates the case each level set of a potential function $f$ is a deformed contact structure. Then the metric can be written in an explicit way and analysis of the metric’s smoothness leads to Theorem D. Finally, the proofs of Theorem A, Corollaries B and C will be given in Section 5 combining our earlier developments.
1.1. Acknowledgment
We benefit greatly from discussion with Profs. Ronan Conlon, McKenzie Wang, Catherine Searle, and Detang Zhou. We are also grateful to anonymous referees for constructive comments and suggestions.
2. Preliminaries
In this section, we recall preliminary results and certain observations which will be used throughout the article.
2.1. GRS and Kählerity
Here we recall the definition of a gradient Ricci soliton, the Kähler setup, and some identities. A Riemannian manifold $(M^{n},g)$ with a potential function $f$ is called a GRS if, for $\mathrm{Rc}$} denoting the Ricci curvature and $\mathcal{L}$ the Lie derivative,
$\text{Rc}+\frac{1}{2}\mathcal{L}_{\nabla f}g={\mathrm{Rc}}+\mathrm{Hess}f=\lambda g.$
We also recall the following well-known identities [27]:
[…]
(2.2) ${\mathrm{Rc}}(\nabla{f})$ $S,\displaystyle=\frac{1}{2}\nabla{\SS}$,
[…]
In the presence of a complex structure, there are further observations. When a manifold $M$ is of an even dimension, an almost complex structure is defined to be a smooth section $J$ of the bundle of endormorphisms $\text{End}(TM)$ such that $^{2}=-\mathrm{Id}. JJ$ is said to be integrable if it is induced from an atlas of complex charts with holomorphic transition functions. $(M,g,J)$ is called an almost Hermitian manifold and $g$ a Hermitian metric if $g(JX,JY)=g(X,Y).$ The fundamental $2$-form or Kähler form is given by $\omega_{g}(X,Y)=g(X,JY).$ $(M,g,J)$ is called almost Kähler if $d\omega=0$. When $J$ is integrable, one upgrades an almost Hermitian to Hermitian and almost Kähler to Kähler. For a Riemannian manifold to be Kähler, the following is well-known.
Proposition 2.1.
[12, Proposition 3.1.9] Let $(M,g,J)$ be an almost Hermitian real manifold. $(M,g,J)$ is Kähler iff $\nabla J=0$
Definition 2.1.
$(M,g,J,f)$ is a Kähler GRS if $(M,g,f)$ is a GRS and $(M,g,J)$ is a Kähler manifold.
It is crucial to observe that, on a Kähler manifold $(M,g,J)$, $\mathrm{Rc}$} is $J$-invariant. Thus, the following is well-known [35, 44, 19].
Lemma 2.2.
Let $(M,g,J)$ be a Kähler manifold and $f:M\mapsto\mathbb{R}$ such that $\mathrm{Hess}f$ is $J$-invariant. Then, we have the following:
  1. (i)
    $J(\nabla f)$ is a Killing vector field.
  2. (ii)
    $\nabla f$ is an infinitesimal automorphism of $J$.
2.2. Foliation
In this subsection, we recall the concept of a foliation and related properties. The references are [67, 12]. A $p$-dimensional foliation $\mathcal{F}$ of an $n$-dimensional manifold $M$ refers to the partition of $M$ into a union of disjoint $p$-dimensional immersed submanifolds $\{L_{\alpha}\}_{\alpha\in A}$, called leaves, with the following property. Every point is on a chart $U$ with coordinates $(x_{1},...x_{p};~y_{1},...y_{q}),~p+q=n,$ such that, for each $L_{\alpha}$, a connected component of $U\cap L_{\alpha}$ is described by the equations
$y_{1}=\text{constant},...,y_{q}=\text{constant}.$
Consequently, it is called a foliated coordinate chart and let $E$ denote the sub-bundle of $TM$ tangential to $\mathcal{F}$. A vector field $X$ is said to be foliate with respect to foliation $\mathcal{F}$ if for every section $Y\in E$, $\mathcal{L}_{X}Y$ is also tangential.
A Riemannian metric $g$ induces an orthogonal decomposition $TM=E\oplus E^{\perp}$. A section of $E^{\perp}$ is said to be horizontal.
Definition 2.2.
A Riemannian metric is said to be bundle-like with respect to a foliation $\mathcal{F}$ if for any foliate horizontal vector fields $X,Y$ and a tangential $V$, $Vg(X,Y)=0.$ In that case, the foliation is said to be Riemannian.
Remark 2.3.
Locally a foliation looks like a submersion and a Riemannian foliation corresponds to a Riemannian submersion.
The following will be important to our investigation.
Proposition 2.3.
[12, Propositions 2.6.7 and 2.6.9] A $1$-dimensional foliation induced by a Killing vector field is Riemannian. Also, a $1$-dimensional Riemannian foliation whose orbits are geodesics is isometric; that is, the flow generated by the associated vector field is an isometry.
For a Riemannian foliation, the toolkit originally developed by B. O’Neill to study submersion [51] will play a crucial role. We let $\pi$ and $\pi^{\perp}$ be the projections from $TM$ onto $E$ and $E^{\perp}$ accordingly. Let $\nabla^{\perp}$ denote the following induced connection, for $Y$ a smooth section of $E^{\perp}$:
[…]
It is verified that $\nabla^{\perp}$ is the unique Levi-Civita connection with respect to the transverse metric $g^{\perp}:=g_{\mid_{E^{\perp}}}$ [12, Exercise 2.4]. Furthermore, we make the simplifying assumption that each leaf of $\mathcal{F}$ is totally geodesic. Let $U$ be a smooth section of $E$ and $X,Y$ be ones of $E^{\perp}$. In this setup, the connection is encoded by tensor $A$ as follows:
$A_{X}U$ $:=\pi^{\perp}(\nabla_{X}U),$
$A_{X}Y$ $:=\pi(\nabla_{X}Y)=-A_{Y}X=\frac{1}{2}\pi([X,Y]).$
2.3. Almost Contact and Contact Structures
In this subsection, we give a brief introduction to almost contact geometry. The reference is [12].
Definition 2.4.
An odd dimensional manifold $P$ is called almost contact if there exists a triple $(\zeta,\eta,\Phi)$ where $\zeta$ is a vector field, $\eta$ is a $1$-form, $\Phi$ is a tensor field of type $(1,1)$, and they satisfy, everywhere on $M$,
$\eta(\zeta)=1\text{ and }\Phi^{2}=-\mathrm{Id}+\zeta\otimes\eta.$
It is immediate from the definition that $\zeta$ is nowhere vanishing and, by Frobenius theorem, $\zeta$ generates an $1$-dimensional foliation $\mathcal{F}$ ($\zeta$ is also called the Reeb vector field [58]). Additionally, $\Phi$ is non-degenerate on the transverse subspace and
$\Phi(\zeta)=0,~~~\eta\circ\Phi=0.$
In the presence of a Riemannian metric $g$, $(P,\zeta,\eta,\Phi)$ is called an almost contact metric structure if $g$ is compatible with $\Phi$. That is,
$g(\Phi(X),\Phi(Y))=g(X,Y)-\eta(X)\eta(Y).$
It is naturally of great interest to consider the transverse geometry $g^{\perp}$, which is the restriction of $g$ to the subspace transverse to $\zeta$:
$g=g^{\perp}+\eta\otimes\eta.$
In our analysis, it is important to determine when leaves of $\mathcal{F}$ are geodesics.
Lemma 2.4.
Let $(P,\zeta,\eta,\Phi)$ be an almost contact metric structure. Then the following are equivalent:
  1. (i)
    Integral curves of $\zeta$ are geodesics.
  2. (ii)
    $\eta$ is invariant along the flow of $\zeta$.
  3. (iii)
    $(\nabla_{\zeta}\Phi)\zeta=0$.
  4. (iv)
    $d\eta(\zeta,\cdot)=0$.
Proof.
It follows from standard calculation; see [12].
Furthermore, the tensor $A$ can be computed immediately [12, Prop 2.5.14].
Lemma 2.5.
Let $(P,\zeta,\eta,\Phi)$ be an almost contact metric structure such that $g$ is bundle-like with respect to the foliation generated by $\zeta$. Then, for horizontal vector fields $X$ and $Y$,
$2g(A_{X}Y,\zeta)$ $=-2d\eta(X,Y)=g([{X},{Y}],\zeta),$
$g(A_{X}\zeta,Y)$ $=d\eta(X,Y).$
The $\Phi$-sectional curvature of an almost contact manifold $(P,\zeta,\eta,\Phi)$ is defined on the horizontal sub-bundle (perpendicular to $\zeta$), for unit length $X$,
$K_{\Phi}(X)=K(X,\Phi(X)).$
Definition 2.5.
An almost contact metric structure $(P,\zeta,\eta,\Phi,g)$ is called contact if one further assumes $g(X,\Phi(Y))=d\eta(X,Y).$
Additionally, a contact metric manifold $(P,\zeta,\eta,\Phi,g)$ is called K-contact if $\zeta$ is Killing; that is, $\mathcal{L}_{\zeta}g=0.$ We are interested in certain $K$-contact structures which give a concrete bridge from almost contact (contact) to almost complex (complex).
Definition 2.6.
A Sasakian structure is a contact metric structure such that the cone $(P\times\mathbb{R}^{+},r^{2}g+dr^{2})$ is Kähler. Moreover, the Kähler form is given by $d(r^{2}\eta)$.
Next we recall a transformation defined in [64].
Definition 2.7.
For $H,F\in\mathbb{R}^{+}$, a $(\pm,H,F)$-deformation is given by
$\zeta^{\ast}=\frac{\zeta}{H},~~~\eta^{\ast}=H\eta,~~~\Phi^{\ast}=\pm\Phi,~~~g^{\ast}=F^{2}g+(H^{2}-F^{2})\eta\otimes\eta.$
A deformed contact/Sasakian metric structure is obtained via an $(\pm,H,F)$-deformation of a contact/Sasakian metric structure.
Remark 2.8.
A $(+,H,F)$ transformation of a contact metric structure is contact iff $H=F^{2}$; this special case is called the transverse homothety deformation [12, Page 228].
It is of interest to relate the curvature of a $(\pm,H,F)$ deformation with one of the original metric. First, there is a simple observation.
Lemma 2.6.
Let $(P,g^{\ast},\zeta^{\ast},\eta^{\ast},\Phi^{\ast})$ be an $(\pm,H,F)$ deformation of an almost contact metric structure. Then we have the following:
  1. (i)
    $(P,g^{\ast},\zeta^{\ast},\eta^{\ast},\Phi^{\ast})$ is an almost contact metric structure.
  2. (ii)
    $g$ is bundle-like with respect to $\mathcal{F}$ if and only if $g^{\ast}$ is.
  3. (iii)
    $\mathcal{L}_{\zeta}g=0\iff\mathcal{L}_{\zeta^{\ast}}g^{\ast}=0$.
Proof.
The proof is via straightforward verification. For example,
$g^{\ast}(\Phi^{\ast}X,\Phi^{\ast}Y)$ $=(F^{2}g+(H^{2}-F^{2})\eta\otimes\eta)(\Phi X,\Phi X)$
$=F^{2}(g(X,Y)-\eta(X)\eta(Y))=g^{\ast}(X,Y)-\eta^{\ast}(X)\eta^{\ast}(Y).$
For $g=g^{\perp}+\eta\otimes\eta$,
${g^{\ast}}=H^{2}\eta\otimes\eta+F^{2}g^{\perp}.$
In case of a K-contact structure, the curvature calculation is relatively simple.
Proposition 2.7.
Let $(P,\zeta,\eta,\Phi,g)$ be a K-contact structure and $(P,g^{\ast},\zeta^{\ast},\eta^{\ast},\Phi^{\ast})$ be its $(\pm,H,F)$ deformation. Then the curvatures are related by, for orthonormal horizontal vectors $X$ and $Y$
${K^{\ast}}(X,Y)$ $=\frac{1}{F^{2}}K^{\perp}(FX,FY)-3\frac{H^{2}}{F^{4}}g^{\ast}(X,\Phi^{\ast}Y)^{2},$
${K^{\ast}}(X,\zeta^{\ast})$ $=\frac{H^{2}}{F^{4}},$
${{\mathrm{Rc}}^{\ast}}(X,Y)$ $={\mathrm{Rc}}^{\perp}(X,Y)-2\frac{H^{2}}{F^{4}}g^{\ast}(X,Y)$
$\ {\mathrm{Rc}}^{\ast}(\zeta^{\ast},\zeta^{\ast})$ $=\frac{H^{2}}{F^{4}}(\text{dim}(M)-1).$
Here $K^{\perp}$ and $\mathrm{Rc}$$^{\perp}$ are the sectional and Ricci curvature of $(g^{\perp},\nabla^{\perp})$.
Proof.
Since $\zeta$ is a Killing vector field, its generated foliation $\mathcal{F}$ is Riemannian by Proposition 2.3. By Lemma 2.6, $(M,g^{\ast},\zeta^{\ast},\eta^{\ast},\Phi^{\ast})$ is an almost contact metric structure, $g^{\ast}$ is bundle-like with respect to $\mathcal{F}$, and $\mathcal{L}_{\zeta}g=0=\mathcal{L}_{\zeta^{\ast}}g^{\ast}$. Therefore, by Lemma 2.4, orbits of $\mathcal{F}$ are geodesics with respect to either $g$ or $g^{\ast}$.
For a Riemannian totally geodesic foliation, the curvature can be computed via tensor AA; see [7, Chapter 9] or [12, Theorem 2.5.16]. In our case, the computation of $A$ is given by Lemma 2.5. ∎
2.4. Rectifiable, Transnormal, and Isoparametric Functions
In this subsection, for a complete connected Riemannian manifold $(M,g)$, we consider a smooth function $f:M^{n}\rightarrow\mathbb{R}$. First, we recall the definition of rectifiable which appears in [56, 22].
Definition 2.9.
$f$ is rectifiable if $|\nabla f|$ is constant along each connected component of level sets of $f$.
A connected component of a regular level set is called a regular connected component.
Lemma 2.8.
The following are equivalent:
  1. (i)
    f is rectifiable.
  2. (ii)
    Integral curves of $\nabla f$ are geodesics after reparametrization.
  3. (iii)
    The gradient of $f$ is an eigenvector of its Hessian.
Proof.
Integral curves of $\nabla f$ are reparametrized geodesics if and only if, when $\nabla f\neq 0$,
$0$ $=\frac{1}{|\nabla f|}\nabla_{\nabla f}\frac{\nabla f}{|\nabla f|}$
$=\frac{1}{|\nabla f|^{3}}\big(|\nabla f|\nabla_{\nabla f}\nabla f-(\nabla_{\nabla f}|\nabla f|)\nabla f\big).$
Thus, it is equivalent to that, for any unit vector field $E_{i}\perp\nabla f$,
$0=g(\nabla_{\nabla f}\nabla f,E_{i})=\mathrm{Hess}f(\nabla f,E_{i})=\frac{1}{2}\nabla_{E_{i}}|\nabla f|^{2}.$
Equivalently, $|\nabla f|^{2}$ is constant on each regular connected component. The equation above also is equivalent to that $\nabla f$ is an eigenvector of $\mathrm{Hess}f$. ∎
Remark 2.10.
Consequently, it justifies the terminology that a rectifiable function is also called one with a geodesic gradient [31, 32].
A priori, $|\nabla f|$ might vary between different connected components. Thus, it is useful to have a global condition.
Definition 2.11.
$f$ is called transnormal if there is a continuous function $b:\mathbb{R}\mapsto\mathbb{R}$ such that $|\nabla f|^{2}=b\circ f.$ Furthermore, a transnormal function is called isoparametric if there is a continuous function $a:\mathbb{R}\mapsto\mathbb{R}$ such that $\Delta f=a\circ f.$
Generally speaking, the former condition corresponds to equidistant level sets while the latter implies that each regular one has constant mean curvature. Also, this subject is closely related to the theory of a transnormal system and polar foliations [48]. It is immediate that if $f$ is rectifiable and each level set is connected then $f$ is transnormal.
Remark 2.12.
A geodesic segment $\gamma:[\alpha,\beta]\mapsto M$ is called an $f$-segment if $\gamma^{\prime}(t)=\frac{\nabla f}{|\nabla f|}$ whenever $|\nabla f|\neq 0$. At a point p such that $\nabla f(p)\neq 0$, it is possible to reparametrize an $f$-segment $\gamma$ via translation such that $\gamma\circ f=\mathrm{Id}$ in a neighborhood of $p$. If $f(p)$ is a local minimum or maximum, one can only reparametrize to obtain such a property in an one-sided neighborhood. Also, with this parametrization, $\frac{\partial f}{\partial t}=|\nabla f|$.
For $c\in f(M)$, denote $M_{c}=f^{-1}(c)$.
Lemma 2.9.
If $f$ is transnormal and $\alpha,\beta]\subset f(M)$ contains no critical value of $f$ then for any $x\in M_{\alpha}$, $y\in M_{\beta}$, we have
  1. (i)
    $d(x,M_{\beta})=d(M_{\alpha},y)=\int_{\alpha}^{\beta}\frac{df}{\sqrt{|\nabla f|}}$;
  2. (ii)
    the integral curves of $\nabla f$ after reparametrization are $f$-segments;
  3. (iii)
    the $f$-segments are the shortest curves among all curves connecting $M_{\alpha}$ and $M_{\beta}$.
Proof.
See [70, Lemma 1]. ∎
Obviously, there is a local version when $f$ is rectifiable. Thus, a Riemannian manifold with a rectifiable function $f$ is locally foliated by equidistant hypersurfaces and one can write the metric in specific way. Following [34, Section 2], let $P$ be a differentiable manifold corresponding to a regular connected component. There is a local diffeomorphism $\phi:I\times P\mapsto M$ such that the metric can be written as $\phi^{\ast}g=dt^{2}+g_{t}.$ Here $t$ is a parametrization of $f(M)$ with unit tangent vector (see Remark 2.12) and $g_{t}$ is an one-parameter family of metrics $P$ such that, for $\phi_{c}:\{c\}\times P\mapsto M_{c}$ the restriction of $\phi$ to a slice, $\phi_{c}^{\ast}g_{\mid M_{c}}=g_{c}.$
For a level set $M_{c}$, we denote the normal exponential map $\Pi_{c}:T^{\perp}M_{c}\mapsto M.$ At a regular value, $\Pi_{c}$ induces a diffeomorphism between nearby regular connected components of distance $\epsilon$: $\Pi_{c}^{\epsilon}:M_{c}\mapsto M_{c+\epsilon}$, via normal geodesics of length $\epsilon$ in the direction of $\partial_{t}$.
Lemma 2.10.
Let $I$ be a continuous open interval with regular values of $f$. For $(c-\epsilon,c+\epsilon)\subset I$, then $\Pi_{c}^{\epsilon}$ is just the identification by the diffeomorphism $\phi$.
Proof.
For $p$ a point in $P$ and $c\in I$ such that $\phi(p,c)=q\in M_{c}$. Let $\gamma(t)$ be the curve given by, for $t\in(c-\epsilon,c+\epsilon)$, $\gamma(t)=\phi(p,t).$ It is readily verified that $\frac{d}{dt}\gamma=\phi_{\ast}\partial_{t}.$ Thus, $\gamma$ is a geodesic segment and by the uniqueness of a geodesic given initial conditions, we conclude that
$\Pi^{\epsilon}_{c}(q)=\gamma(\epsilon)=\phi(p,\epsilon).$
A soliton structure comes with a potential function and Lemma 2.8 is refined as follows, which is also implicit in [56].
Proposition 2.11.
Let $(M,g,f)$ be a GRS with a non-constant $f$. The following are equivalent:
  1. (i)
    f is rectifiable,
  2. (ii)
    whenever $\nabla f\neq 0$, $\nabla f$ is an eigenvector of $\mathrm{Rc}$},
  3. (iii)
    $S\nabla f\parallel\nabla\SS$ everywhere.
Furthermore, $f$ is transnormal iff $f$ is isoparametric.
Proof.
$f$ is rectifiable iff $|\nabla f|$ is constant on each regular connected component and, by equation (2.3), so is $S\SS$. By Sard’s theorem, the set of regular values is dense. Thus, by continuity, it is equivalent to $S\nabla f\parallel\nabla\SS$ everywhere. That establishes $(i)\iff(iii)$. Finally, $(ii)\iff(iii)$ follows from equation (2.2).
When $f$ is rectifiable, $b:f(M)\mapsto{\mathrm{R}}$ can be defined locally and is smooth at a regular value. At a critical one, it is only smooth in the sense of one-sided limits. Furthermore, the local foliation of equidistant hypersurfaces allows one to rewrite the soliton equation as follows [30]. For $N=\partial_{t}$, the shape operator is given as $LX:=\nabla_{X}N.$ Denoting the ordinary derivative $\frac{d}{dt}$ by , it follows that
(2.5) $g^{\prime}$ $=2g\circ L,~~~~~~\nabla_{N}L=L^{\prime}.$
Due to Gauss, Codazzi, and Riccati equations, the Ricci curvature of $(M,g)$ is determined by that of $(M_{t},g_{t})$ and the shape operator. That is,
${\mathrm{Rc}}(X,Y)$ $={\mathrm{Rc}}_{t}(X,Y)-\text{tr}(L_{t})g_{t}(LX,Y)-g_{t}({L^{\prime}}(X),Y),$
(2.6) ${\mathrm{Rc}}(X,N)$ $=-\nabla_{X}\text{tr}(L_{t})-g_{t}(\delta L,X),$
${\mathrm{Rc}}(N,N)$ $=-\text{tr}({L^{\prime}})-\text{tr}(L^{2}).$
Here ${\mathrm{Rc}}_{t}$ denotes the Ricci curvature of $(M_{t},g_{t})$, $\text{tr}{T}=\text{tr}_{g_{t}}T_{t}$, and $\delta$ the co-differential. Next, we recall $\mathrm{Hess}f(X,Y)=g(\nabla_{X}{\nabla f},Y)$ and $\nabla f=\frac{df}{dt}N=f^{\prime}N$. Consequently, the gradient Ricci soliton equation ${\mathrm{Rc}}+\mathrm{Hess}{f}=\lambda g$ is reduced to:
$0$ $=-(\delta L)-\nabla\text{tr}{L},$
(2.7) $\lambda$ $=-\text{tr}({L^{\prime}})-\text{tr}(L^{2})+f^{\prime\prime},$
$\lambda g(X,Y)$ $={\mathrm{Rc}}(X,Y)-(\text{tr}{L})g(LX,Y)-g({L^{\prime}}(X),Y)+f^{\prime}g(LX,Y).$
3. Deformed Sasakian
In this section, we give a characterization for an almost contact metric manifold to be a deformed Sasakian structure, proving Theorem E. Consequently, it is possible to detect such a structure on a real hypersurface of a Kähler manifold via examining its shape operator as in Theorem F. Thus, they generalize classical results, [12, Theorem 7.3.16 and Theorem 7.3.18], and might be of independent interest.
Let $(P,\zeta,\eta,\Phi,g)$ be an almost contact metric manifold and $\mathcal{F}$ be the foliation generated by $\zeta$. The following is well-known.
Theorem 3.1.
[12, Theorem 7.3.19] $(P,\zeta,\eta,\Phi,g)$ is Sasakian if and only if, for every vector fields $X$ and $Y$,
(3.1) $\big(\nabla_{X}\Phi\big)Y=g(X,Y)\zeta-\eta(Y)X.$
Towards a generalization, we observe the following.
Lemma 3.2.
If $\big(\nabla_{X}\Phi\big)Y=bg(X,Y)\zeta-b\eta(Y)X$ for a real number $b$ and any vector fields $X$ and $Y$ then $g$ is bundle-like with respect to $\mathcal{F}$ and $\zeta$ is Killing.
Proof.
The assumption implies that, interchanging $X$ and $Y$,
$b(\eta(X)Y-\eta(Y)X)$ $=\big(\nabla_{X}\Phi\big)Y-\big(\nabla_{Y}\Phi\big)X=\nabla_{X}(\Phi Y)-\nabla_{Y}(\Phi X)-\Phi([X,Y])$
Let $X=\zeta$ and $Y$ be a foliate horizontal vector field then $[X,Y]$ is tangential and, thus,
$\nabla_{Y}(\Phi X)=\Phi([X,Y])=0.$
Consequently, $\nabla_{\zeta}(\Phi Y)=bY.$ Applying the assumption again yields
$0$ $=bg(\zeta,\Phi Y)\zeta-\eta(\Phi(Y))b\zeta=\nabla_{\zeta}\Phi(\Phi Y)$
$=\nabla_{\zeta}\Phi^{2}(Y)-\Phi(\nabla_{\zeta}(\Phi Y))=\nabla_{\zeta}(-Y)-b\Phi(Y).$
Thus, for foliate horizontal vector fields $Y$ and $Z$,
$\zeta g(Y,Z)$ $=g(\nabla_{\zeta}Y,Z)+g(Y,\nabla_{\zeta}Z)=-bg(\Phi Y,Z)-bg(Y,\Phi Z)=0.$
The last equality is due to the compatibility of $g$ and $\Phi$. Therefore, $g$ is bundle-like by Definition 2.2 and $\mathcal{F}$ is a Riemannian foliation. Furthermore, Lemma 2.4 is also applicable since $\nabla_{\zeta}\Phi(\zeta)=b\zeta-b\zeta=0.$ Thus $\mathcal{F}$ is a Riemannian foliation whose orbits are geodesics. By Proposition 2.3, $\zeta$ is a Killing vector field.
Lemma 3.3.
For a horizontal vector field $X$, the following are equivalent
  1. (i)
    $(\nabla_{\zeta}\Phi)X=0$
  2. (ii)
    $\zeta,\Phi(X)]-\Phi([\zeta,X])=\Phi(A_{X}\zeta)-A_{\Phi X}\zeta$.
Proof.
We compute, using the notation from Section 2,
$(\nabla_{\zeta}\Phi)(X)$ $=\nabla_{\zeta}(\Phi X)-\Phi(\nabla_{\zeta}X)$
$=[\zeta,\Phi X]-\Phi([\zeta,X])-\Phi(\nabla_{X}\zeta)+\nabla_{\Phi X}\zeta,$
$=[\zeta,\Phi X]-\Phi([\zeta,X])-\Phi(A_{X}\zeta)+A_{\Phi X}\zeta.$
Moreover, since $\mathcal{F}$ is Riemannian, by Lemma 2.5, the tensor $A$ can be computed as, for horizontal vector fields $X$ and $Y$,
$2A_{X}Y$ $=-2d\eta(X,Y)\zeta=g([{X},{Y}],\zeta)\zeta$
$g(A_{X}\zeta,Y)$ $=d\eta(X,Y).$
Lemma 3.4.
For horizontal vector fields $X,Y$, the following are equivalent
  1. (i)
    $(\nabla_{X}\Phi)Y=bg(X,Y)\zeta$
  2. (ii)
    $A_{X}(\Phi Y)=bg^{\perp}(X,Y)\zeta$ and $\nabla_{X}^{\perp}(\Phi Y)=\Phi(\nabla_{X}^{\perp}Y)$
Proof.
We compute,
$(\nabla_{X}\Phi)(Y)$ $=(\nabla_{X}(\Phi Y)-\Phi(\nabla_{X}Y)$
$=A_{X}(\Phi Y)+\pi^{\perp}(\nabla_{X}(\Phi Y))-\Phi(A_{X}Y+\pi^{\perp}(\nabla_{X}Y))$
$=A_{X}(\Phi Y)+\nabla^{\perp}_{X}(\Phi Y))-\Phi(\nabla^{\perp}_{X}Y).$
Lemma 3.5.
For horizontal vector fields $X,Y$, the following are equivalent
  1. (i)
    $(\nabla_{X}\Phi)\zeta=-bX$
  2. (ii)
    $A_{X}\zeta=-b\Phi X$.
Proof.
We compute
$(\nabla_{X}\Phi)(\zeta)$ $=(\nabla_{X}(\Phi(\zeta))-\Phi(\nabla_{X}\zeta)$
$=-\Phi(A_{X}\zeta).$
Thus, it is possible to track how the covariant derivative of $\Phi$ changes under a deformation.
Theorem 3.6.
Let $(P,\zeta,\eta,\Phi,g)$ be an almost contact metric structure such that, for every vector fields $X$ and $Y$,
$\nabla_{X}\Phi(Y)=bg(X,Y)\zeta-\eta(Y)bX.$
Let $(P,\zeta^{\prime},\eta^{\prime},\Phi^{\prime},g^{\prime})$ be an $(\pm,1,c)$-deformation then
$\nabla^{\prime}_{X}\Phi^{\prime}(Y)=\pm\frac{b}{c^{2}}\big(g(X,Y)\zeta^{\prime}-\eta(Y)X\big).$
Proof.
Let $\mathcal{F}$ be the foliation generated by $\zeta$. By Lemma 3.2, $\mathcal{F}$ is Riemannian and $\zeta$ is a Killing vector field. We write the metric as
$g=g^{\perp}+\eta\otimes\eta.$
By Lemmas 3.4 and 3.5, for horizontal vector fields $X,Y$,
$A_{X}\Phi(Y)$ $=bg(X,Y)\zeta,$
(3.2) $\nabla^{\perp}_{X}\Phi(Y)$ $=\Phi(\nabla_{X}^{\perp}Y),$
$A_{X}\zeta$ $=-b\Phi(X).$
We will assume $(P,\zeta^{\prime},\eta^{\prime},\Phi^{\prime},g^{\prime})$ be an $(1,c)$-deformation as the minus case can be done similarly. Thus, $\eta^{\prime}=\eta$, $\zeta^{\prime}=\zeta$, $\Phi^{\prime}=\Phi$ and
$g^{\prime}=c^{2}g+(1-c^{2})\eta\otimes\eta=c^{2}g^{\perp}+\eta\otimes\eta.$
By Lemma 2.6, $(P,g^{\prime},\eta^{\prime},\zeta^{\prime},\Phi^{\prime})$ is an almost contact metric structure, $\zeta=\zeta^{\prime}$ is a Killing unit vector field with respect to $g^{\prime}$, and $g^{\prime}$ is bundle-like with respect to $\mathcal{F}$. By Lemma 2.5 and equation (3.2), we have
$A^{\prime}_{X}(\Phi^{\prime}Y)$ $=-d\eta^{\prime}(X,\Phi^{\prime}(Y))\zeta^{\prime}=d\eta(X,\Phi(Y))\zeta=A_{X}\Phi(Y)$
$=bg^{\perp}(X,Y)\zeta=\frac{b}{c^{2}}g^{\prime}(X,Y)\zeta^{\prime}.$
Similarly,
$g^{\prime}(A^{\prime}_{X}\zeta^{\prime},Y)$ $=d\eta^{\prime}(X,Y)=d\eta(X,Y)=g(A_{X}\zeta,Y)$
$=-g(b\Phi(X),Y)=-\frac{b}{c^{2}}g^{\prime}(\Phi(X),Y).$
Thus, $A^{\prime}_{X}\zeta=-\frac{b}{c^{2}}\Phi(X)$. As the Levi-Civita connection is scaling invariant, $\nabla^{c^{2}g^{\perp}}=\nabla^{g^{\perp}}=\nabla^{\perp}$. By Lemma 3.3, $\nabla_{\zeta}\Phi(X)=0$ if and only if
$[\zeta,\Phi(X)]-\Phi([\zeta,X])$ $=\Phi(A_{X}\zeta)-A_{\Phi X}\zeta=\Phi(-b\Phi(X))+b\Phi(\Phi(X))=0.$
As the left-hand side is independent of the metric, $\nabla^{\prime}_{\zeta^{\prime}}\Phi^{\prime}(X)=(\nabla_{\zeta}\Phi)(\zeta)=0$. Together with Lemmas 3.4 and 3.5, we obtain, for any vector fields $V$ and $W$,
$\nabla^{\prime}_{V}\Phi^{\prime}(W)=\frac{b}{c^{2}}\Big(g^{\prime}(V,W)\zeta^{\prime}-\eta^{\prime}(W)V\Big).$
Corollary 3.7.
Let $(P,\zeta,\eta,\Phi,g)$ be a Sasakian manifold and $(P,\zeta^{\prime},\eta^{\prime},\Phi^{\prime},g^{\prime})$ be a $(\pm,H,F)$-deformation. Then
$\nabla^{\prime}_{X}\Phi^{\prime}(Y)=\frac{\pm H}{F^{2}}\big(g(X,Y)\zeta^{\prime}-\eta(Y)X\big).$
Proof.
By Remark 2.8, a $(H,\sqrt{H})$-deformation is Sasakian. Thus, applying Theorem 3.6 for $b=1,c=\frac{F}{\sqrt{|H|}}$ in combination with Theorem 3.1 yields the result. ∎
We are ready to give the proof of Theorem E.
Proof of Theorem E.
For part (i), without loss of generality, we assume $b>0$ (if $b<0$ then we consider $(P,\zeta,\eta,-\Phi$). One direction follows from Corollary 3.7. The other is deduced by applying Theorem 3.6 for $c=\sqrt{b}.$
For part (ii), $b=0$ means $\Phi$ is parallel since, for every vector fields $X$ and $Y$
$\nabla_{X}\Phi(Y)=0.$
By Lemmas 2.4 and 3.2, $\zeta$ is a Killing and geodesic vector field. By Lemma 3.5, for any horizontal vector field $X$,
$\nabla_{X}\zeta=A_{X}\zeta=0.$
Thus, $\zeta$ is a unit-length parallel vector field and, accordingly, its universal cover splits into a Riemannian product. The factor perpendicular to $\zeta$ has $\Phi$ as an almost complex structure. Its compatibility with $g$ makes the induced metric almost Herminian. Since $\Phi$ is parallel, the metric is Kähler by Proposition 2.1. ∎
Proof of Theorem F.
Let $\nabla$ and $\nabla^{P}$ denote the Levi-Civita connection on $(M,g)$ and $P$ with the induced metric, respectively. We have
$\nabla_{X}Y=(\nabla^{P}_{X}Y)-g(LX,Y)V.$
Since $J$ is parallel with respect to $\nabla$, $\nabla_{X}(JY)=J(\nabla_{X}Y)$. Thus,
$\nabla^{P}_{X}(\Phi Y)-g(LX,\Phi Y)V+X\eta(Y)V+\eta(Y)LX$ $=\Phi(\nabla^{P}_{X}Y)+\eta(\nabla_{X}Y)V+g(LX,Y)\zeta.$
Consequently, equating the normal and tangent components give
$(\nabla^{P}_{X}\Phi)Y$ $=-\eta(Y)LX+g(LX,Y)\zeta,$
$-g(LX,\Phi Y)+X\eta(Y)$ $=\eta(\nabla_{X}Y)$
For the first equation, substituting $L=\alpha\mathrm{Id}+\beta\zeta\otimes\eta$ yields
$(\nabla^{P}_{X}\Phi)Y$ $=-\eta(Y)\Big(\alpha X+\beta\zeta\eta(X)\Big)+g\Big(\alpha X+\beta\zeta\eta(X),Y\Big)\zeta$
$=-\alpha\eta(Y)X+\alpha g(X,Y)\zeta.$
Similarly, for the second equation, we have
$\eta(\nabla_{X}Y)$ $=-g(\alpha X+\beta\zeta\eta(X),\Phi Y)+X\eta(Y),$
$\iff g(\nabla_{X}\zeta,Y)$ $=\alpha g(X,\Phi(Y))=\alpha g(\Phi(X),-Y),$
$\iff\nabla_{X}\zeta$ $=-\alpha\Phi(X).$
By Lemmas 3.2 and 3.5, the second equation is implied by the first. The result then follows from Theorem E. ∎
4. Kähler Solitons and Contact Structures
In this section, we’ll study a Kähler GRS $(M,g,J,f)$ under the assumption that each regular connected component level set of $f$ is a deformed contact structure. Thus, a proof of Theorem D will be provided. For $c\in f(M)$, we recall the construction (1.2), for a regular value $c$,
$V:=\frac{\nabla f}{|\nabla f|},~~g_{c}:=g_{\mid M_{c}},~~\zeta_{c}:=-J(V),~~\eta_{c}(\cdot):=g(\cdot,\zeta_{c}),~~\Phi_{c}(\cdot):=-\eta_{c}(\cdot)V+J(\cdot).$
Together, $(M_{c},\zeta_{c},\eta_{c},\Phi_{c},g_{c})$ is an almost contact metric structure. Additionally, the shape operator is given by
$g(LX,Y)=g(\nabla_{X}V,Y)=g\big(\nabla_{X}\frac{\nabla f}{|\nabla f|},Y\big)=\frac{|\nabla f|\mathrm{Hess}f(X,Y)-X|\nabla f|g(\nabla f,Y)}{|\nabla f|^{2}}.$
Thus,
(4.1) $\mathrm{Hess}f(X,Y)=|\nabla f|g(LX,Y)+\frac{X|\nabla f|}{|\nabla f|}g(\nabla f,Y).$
We observe the following.
Lemma 4.1.
If $(M_{c},\zeta_{c},\eta_{c},\Phi_{c},g_{c})$ is a deformed connected contact metric structure, then it is a deformed $K$-contact structure and $|\nabla f|$ is constant along $M_{c}$.
Proof.
For convenience, we drop the $c$ dependence. By Lemma 2.2, $W:=J(\nabla f)=-|\nabla f|\zeta$ is a Killing vector field. It suffices to show $|\nabla f|$ is constant along $M_{c}$. First, we observe that, since $\mathrm{Hess}f$ is $J$-invariant:
$\nabla_{W}|\nabla f|^{2}$ $=2g(\nabla_{W}\nabla f,\nabla f)=2\mathrm{Hess}f(W,\nabla f)=0.$
Next, the deformed contact metric structure implies that
$d\eta(\zeta,\cdot)=0.$
By Lemma 2.4, the integral curves of $\zeta$ are geodesics. Therefore, for any $X$ tangential to $M_{c}$,
$0=g(\nabla_{\frac{W}{|W|}}\frac{W}{|W|},X)$ $=\frac{|W|g(\nabla_{W}W,X)-W|\nabla f|g(W,X)}{|\nabla f|^{3}}.$
Since $W|\nabla f|^{2}=0$ we have $g(\nabla_{W}W,X)=0$. Since $W$ is a Killing vector field,
$0=g(\nabla_{X}W,W)=\frac{1}{2}X|W|^{2}.$
As $|W|=|\nabla f|$ the result follows.
If $f$ is rectifiable, for horizontal vector fields $X,Y$,
$2d\eta_{c}(X,Y)$ $=(\nabla_{X}\zeta_{c},Y)-(\nabla_{Y}\zeta_{c},X),$
$=\frac{1}{|\nabla f|}\Big((\nabla_{X}(-J\nabla f),Y)-(\nabla_{Y}(-J\nabla f),X)\Big),$
(4.2) $=\frac{2}{|\nabla f|}\mathrm{Hess}f(X,JY).$
Here we also use that $\mathrm{Hess}f$ is compatible with $J$.
Lemma 4.2.
Let $(P,\zeta,\eta,\Phi,\eta\otimes\eta+g^{\perp})$ be a contact metric structure. Suppose the metric and the almost complex structure, on $I\times P$,
$g$ $=dt^{2}+H^{2}(t)\eta\otimes\eta+F^{2}(t)g^{\perp},$
$J$ $=\partial_{t}\otimes H\eta-\frac{1}{H}\zeta\otimes dt+\Phi,$
are almost Kähler. Then,
$FF^{\prime}=H.$
Proof.
The contact structure implies $d\eta=g^{\perp}(\cdot,\Phi(\cdot)):=\omega^{\perp}(\cdot,\cdot).$ Recall $\omega_{g}(X,Y)=g(X,JY)$, the Kähler form becomes,
$\omega_{g}$ $=2dt\wedge H\eta+F^{2}\omega^{\perp}.$
Thus, $d\omega_{g}=-2Hdt\wedge\omega^{\perp}+2FF^{\prime}dt\wedge\omega^{\perp}$ and one deduces
$FF^{\prime}=H.$
We are ready to give the main proof of this section.
Proof of Theorem D.
By Lemma 4.1, $f$ is rectifiable and each level set is endowed with a deformed $K$-contact structure. Let $P$ be a differentiable manifold corresponding to a regular connected component of $f$. As in Subsection 2.4, there is a local diffeomorphism $\phi:I\times P\mapsto M$ such that
$\phi^{\ast}(g)=dt^{2}+g_{t}.$
for $(P,g_{t})$ an one-parameter family of Riemannian metrics which are equal to the pullback of induced metric on nearby connected components. We observe, since $f$ is invariant by the Killing vector field $J(\nabla f)$
$0$ $=\mathcal{L}_{J\nabla f}\nabla f=-\mathcal{L}_{\nabla f}J(\nabla f).$
Therefore, $J(\nabla f)=-|\nabla f|_{t}\zeta_{t}$ is invariant by the flow generated by $\nabla f$. Thus, by Lemma 2.10, $f^{\prime}\zeta_{t}$ is identified with a fixed vector $\zeta$ on $P$. Furthermore, equations (2.5) and (4.1) and that $\nabla f$ an eigenvector of $\mathrm{Hess}f$ imply that the transverse subspace is also invariant as $t$ varies. Thus, we can fix a background one-form $\eta$ and write
$g_{t}$ $=H^{2}(t)\eta\otimes\eta+g^{\perp}_{t},$
$H$ $=f^{\prime},$
with $g^{\perp}_{t}$ being the restriction of $g_{t}$ to the transverse subspace. Since each is a deformed contact structure, restricted to the transverse subspace, $d(H\eta)(\cdot,\cdot)$ is a multiple of $g^{\perp}_{t}(J\cdot,\cdot)$. By equation (4.2), $g_{t}^{\perp}$ is a multiple of $\mathrm{Hess}f$ restricted to the transverse subspace. Combining with equations (2.5) and (4.1), one deduces that $\frac{d}{dt}g_{t}^{\perp}$ is a multiple of $g_{t}^{\perp}$ with the scalar only depending on $t$. Thus, for a fixed transverse metric $g^{\perp}$,
$g=dt^{2}+H^{2}(t)\eta\otimes\eta+F^{2}(t)g^{\perp}.$
In particular, nearby regular level sets are endowed with the deformation of the same $K$-contact structure.
Next, using (2.5), we compute the shape operator, for $\mathrm{Id}$ denoting the identity on the transverse subspace and $m=\text{dim}_{\mathbb{C}}M-1$,
$L_{t}$ $=\frac{H^{\prime}}{H}\zeta\otimes\eta+\frac{F^{\prime}}{F}\mathrm{Id}.\,$
${L^{\prime}}_{t}$ $=\Big(\frac{H^{\prime\prime}}{H}-\big(\frac{H^{\prime}}{H}\big)^{2}\Big)\zeta\otimes{\eta}+\Big(\frac{F^{\prime\prime}}{F}-\big(\frac{F^{\prime}}{F}\big)^{2}\Big)\mathrm{Id};$
$\text{tr}{L_{t}}$ $=\frac{H^{\prime}}{H}+(2m)\frac{F^{\prime}}{F},$
$\text{tr}{L^{2}_{t}}$ $=\big(\frac{H^{\prime}}{H}\big)^{2}+(2m)\frac{(F^{\prime})^{2}}{F^{2}},$
$\text{tr}{L^{\prime}_{t}}$ $=\frac{H^{\prime\prime}}{H}+(2m)\frac{F^{\prime\prime}}{F}-\frac{(H^{\prime})^{2}}{H^{2}}-(2m)\frac{(F^{\prime})^{2}}{F^{2}}.$
By Theorem E, at regular values, $F^{\prime}\neq 0$ and $(P,\eta_{t},\zeta_{t},\Phi_{t},g_{t})$ is a deformed Sasakian structure. Combining (2.7) and Proposition 2.7, we have
${\mathrm{Rc}}^{\perp}-2\frac{H^{2}}{F^{4}}g_{t}$ $=\lambda g_{t}+\text{tr}(L)g_{t}\circ L-g_{t}\circ L^{\prime}+f^{\prime}g_{t}\circ L_{t}.$
Since $g_{t}=F^{2}g^{\perp}+H^{2}\eta\otimes\eta$, one deduces that ${\mathrm{Rc}}^{\perp}=kg^{\perp}$. Consequently, the soliton equation (1.1) becomes
$\lambda$ $=-\frac{H^{\prime\prime}}{H}-(2m)\frac{F^{\prime\prime}}{F}+f^{\prime\prime}$
$=\frac{H^{2}}{F^{4}}(2m)-\frac{H^{\prime\prime}}{H}-2m\frac{H^{\prime}F^{\prime}}{HF}+f^{\prime}\frac{H^{\prime}}{H}$
(4.3) $=\frac{k}{F^{2}}-\frac{H^{2}}{F^{4}}2-\frac{F^{\prime\prime}}{F}-(2m-1)\big(\frac{F^{\prime}}{F}\big)^{2}-\frac{H^{\prime}F^{\prime}}{FH}+f^{\prime}\frac{F^{\prime}}{F}.$
By Lemma 4.2, the ODE system is augmented with additional constraints $FF^{\prime}=H=f^{\prime}$. It was investigated and solved explicitly via a transformation to a modified Calabi’s ansatz in [30] (see also [68]). In particular, $f$ is monotonic and one concludes that $f$ is transnormal and each level set is connected.
Furthermore, there must be a finite $t$ where either $H\rightarrow 0$ or both $H,F\rightarrow 0$. By Lemmas 4.3 and 4.4 below, there is a Riemannian submersion from $(P,\eta,\zeta,\Phi,g^{\perp}+\eta\otimes\eta)$ to a Kähler-Einstein manifold. Thus, the result follows.
For the following results, one assumes the setup as in the proof of Theorem D.
Lemma 4.3.
Let $I=(0,\epsilon)$ for $\epsilon>0$ and suppose that
$\lim_{t\rightarrow 0^{+}}H(t)=\lim_{t\rightarrow 0^{+}}F(t)=0,$
and the metric can be extended smoothly to $t=0$. Then, $(P,\zeta,\eta,\Phi,g^{\perp}+\eta\otimes\eta)$ is a Sasakian sphere with Hopf fibration over a complex projective space.
The proof models after one for [55, Theorem 4.3.3]. The idea is that, as $H,F\rightarrow 0$, the metric must become rounder and rounder. The rigidity of the deformed structure implies that it is actually round.
Proof.
Let $p$ denotes the point compactification at $t=0$. Since it is locally Euclidean around $p$, each level set of $f$ corresponds to a distance sphere. This also follows from the Morse lemma as
$\mathrm{Hess}f(p)=\lim_{t\rightarrow 0}f^{\prime\prime}(t)g.$
Thus, $P$ is diffeomorphic to a sphere. Consequently, there is a diffeomorphism $\hat{\phi}$ from an open ball in $\mathbb{R}^{2n}$ to $M$ such that each Euclidean round sphere is mapped to a level set of $f$.
Let’s fixed a sphere with the standard round metric $(\mathbb{S}^{2n-1},g_{round})$. In $\mathbb{R}^{2n}$, the intersection of a $2$-dimensional plane through the origin with $(\mathbb{S}^{2n-1},g_{round})$ is a great circle $\mathbb{S}^{1}$ with coordinate $\theta$. That is,
$g_{round}(\partial_{\theta},\partial_{\theta})=1.$
Next, we consider the image of that plane via $\hat{\phi}$. Since $\hat{\phi}$ is a diffeomorphism, the image is a submanifold of dimension two with coordinates $t$ and $\theta$. Using polar coordinates
$x$ $=t\cos\theta,$
$y$ $=t\sin{\theta},$
$\partial_{x}$ $=\cos{\theta}\partial_{t}-\frac{1}{t}\sin\theta\partial_{\theta},$
$\partial_{y}$ $=\sin{\theta}\partial_{t}+\frac{1}{t}\cos\theta\partial_{\theta}$
Thus, for $g=dt^{2}+H^{2}(t)\eta\otimes\eta+F^{2}(t)g^{\perp}$
$g_{xx}:=g(\partial_{x},\partial_{x})$ $=\cos^{2}\theta+\frac{1}{t^{2}}\sin^{2}{\theta}(H^{2}\eta^{2}(\partial_{\theta})+F^{2}g^{\perp}(\partial_{\theta},\partial_{\theta})),$
$g_{yy}:=g(\partial_{y},\partial_{y})$ $=\sin^{2}\theta+\frac{1}{t^{2}}\cos^{2}{\theta}(H^{2}\eta^{2}(\partial_{\theta})+F^{2}g^{\perp}(\partial_{\theta},\partial_{\theta})).$
Letting $t\rightarrow 0$ yields
$g_{xx}(p)$ $=\cos^{2}\theta+\sin^{2}{\theta}\big((H^{\prime}(0))^{2}\eta^{2}(\partial_{\theta})+(F^{\prime}(0))^{2}g^{\perp}(\partial_{\theta},\partial_{\theta})\big),$
$g_{yy}(p)$ $=\sin^{2}\theta+\cos^{2}{\theta}\big((H^{\prime}(0))^{2}\eta^{2}(\partial_{\theta})+(F^{\prime}(0))^{2}g^{\perp}(\partial_{\theta},\partial_{\theta})\big).$
Adding them together we deduce that, since $g(p)$ is independent of $\theta$,
$1=(H^{\prime}(0))^{2}\eta^{2}(\partial_{\theta})+(F^{\prime}(0))^{2}g^{\perp}(\partial_{\theta},\partial_{\theta})=g_{round}(\partial_{\theta},\partial_{\theta}).$
Since $\partial_{\theta}$ is arbitrary,
$H^{\prime}(0)^{2}\eta\otimes\eta+F^{\prime}(0)^{2}g^{\perp}=g_{round}.$
By Proposition 2.7, $g^{\perp}+\eta\otimes\eta$ has constant $\Phi$-holomorphic sectional curvature. By the classification of Tanno [65], it must be the standard round sphere with the contact structure given by the Hopf fibration over a complex projective space.
Lemma 4.4.
Suppose that
$\lim_{t\rightarrow 0^{+}}H(t)=0\text{ and }\lim_{t\rightarrow 0^{+}}F(t)\neq 0,$
and the metric can be extended smoothly to $t=0$. Then, $(P,\eta,\zeta,\Phi,g^{\perp}+\eta\otimes\eta)$ is a Riemannian submersion over a Kähler-Einstein manifold.
Proof.
By the ODE system (4.3), the set of points corresponding to $t=0$ and $H\rightarrow 0$ corresponds to a level set, called $N$, of a critical value of $f$. NN is also the zero set of the Killing vector field $J(\nabla f)$ and, thus, by Kobayashi’s [43], is a totally geodesic submanifold of an even codimension with induced metric $g_{N}$.
Let’s recall the diffeomorphism between nearby regular level sets, induced by the normal exponential map, $\Pi_{c}^{\epsilon}:M_{c}\mapsto M_{c+\epsilon}$. It will induce an onto submersion $\Pi^{-\epsilon}_{\epsilon}:M_{\epsilon}\mapsto N=M_{0}$ for sufficiently small $\epsilon$. Indeed, for a point $p\in N$, its pre-image via $\Pi^{-\epsilon}_{\epsilon}$ is the image of the exponential map at $p$ of the sphere of radius $\epsilon$ in its normal bundle [70, Lemma 9].
Let $X$ be a vector field on $P$ and its identification on $M_{\epsilon}$ via $\phi$. Let $\widetilde{X}$ be the push-forward of $X$ via $\Pi^{-\epsilon}_{\epsilon}$. By continuity and Lemma 2.10,
[…]
$=F(0)^{2}g^{\perp}(X,X).$
Thus, $\widetilde{X}=\vec{0}$ if and only if $X$ is a multiple of $\zeta$. Thus, at each point $q\in M_{\epsilon}$, the differential of $\Pi^{-\epsilon}_{\epsilon}$ is onto with an one-dimensional kernel. Thus, $N$ is of co-dimension two and $\Pi^{-\epsilon}_{\epsilon}$ is a conformal submersion.
It remains to show $(N,F(0)^{2}g_{N})$ is Kähler-Einstein with a compatible almost complex structure. Indeed, $\Phi$ induces an almost complex structure on $N$ by
$J_{N}\widetilde{X}:=\widetilde{\Phi X}$.
[…]
Proof. Without loss of generality, one assumes that $X\perp\zeta$ and, by the construction of $\Phi$, $\widetilde{\Phi X}=\widetilde{JX}$. On the other hand, $\widetilde{X}=\lim_{t\rightarrow 0}(\phi_{t})_{\ast}X$ and, by continuity,
[…]
$=\lim_{t\rightarrow 0}(\phi_{t})_{\ast}(\phi_{t}^{\ast}JX)=\lim_{t\rightarrow 0}(\phi_{c})_{\ast}(JX)$
$=\widetilde{JX}$.
Claim: $(N,g_{N},J_{N})$ is a Kähler-Einstein manifold.
Proof. By continuity,
[…]
Thus, it is Einstein. Then $J_{N}$ is parallel since $_{N}=J_{\mid TN}, JJ$ is parallel, and $(N,g_{N})$ is totally geodesic.
5. Dimension four
We restrict our investigation to real dimension four and great simplification occurs.
Lemma 5.1.
Let $(M,g,J,f)$ be a Kähler GRS in real dimension four. Suppose that $f$ is rectifiable then along each regular connected component, $\mathrm{Hess}(f)$ has two constant eigenvalues, each of multiplicity two.
Proof.
Let $M_{c}$ be a regular connected component of $f$. Since $|\nabla f|$ is constant on $M_{c}$, for any vector field $X$ tangent to $M_{c}$,
$\text{Hess (f)}(\nabla f,X)=0.$
Following Proposition 2.11, $b(f)=|\nabla f|^{2}$ is locally defined for nearby connected components and it is smooth. Thus,
$\text{Hess (f)}(\nabla f,\nabla f)$ $=g(\nabla_{\nabla f}(\nabla f),\nabla f)$
$=\frac{1}{2}\nabla_{\nabla f}|\nabla f|^{2}=\frac{1}{2}b^{\prime}(c)|\nabla f|^{2}.$
Therefore, $\nabla f$ is an eigenvector with eigenvalue $\frac{1}{2}b^{\prime}(c)$. As each tangent space is of dimension $4$ and Hess f is $J$-invariant, Hess f has exactly two eigenvalues each of multiplicity two. By equations (2.1) and (2.3),
$\Delta f=\text{tr}(\mathrm{Hess}(f))=n\lambda-|\nabla f|^{2}+2\lambda f-\text{a constant}.$
Thus, it is also fixed along $M_{c}$ and the result then follows.
We are now ready to classify all Kähler GRS with a rectifiable potential function $f$.
Proof of Theorem A.
Let $M_{c}$ be a regular connected component of $f$. Then, $(M_{c},\zeta_{c},\eta_{c},\Phi_{c},g_{c})$ via (1.2) is an almost contact metric structure. Along a regular connected component, $|\nabla f|$ is a nonzero constant and we have, for $L$ denoting the shape operator,
$\mathrm{Hess}(f)(X,Y)$ $=|\nabla f|g(LX,Y)+\frac{X|\nabla f|}{|\nabla f|}g(\nabla f,Y),$
$=|\nabla f|g(LX,Y).$
By Lemma 5.1, $\mathrm{Hess}(f)$ has constant eigenvalues $\mu_{1},\mu_{2}$ with $\zeta$ as one of the eigenvector. Therefore,
$L=\frac{1}{|\nabla f|}((\mu_{1}-\mu_{2})\zeta\otimes\eta+\mu_{2}\mathrm{Id}).$
Consequently, $L$ satisfies the condition of Theorem F and $(M_{c},\zeta_{c},\eta_{c},\Phi_{c},g_{c})$ must be either a deformed Sasakian structure or locally a Riemannian product. By continuity, nearby regular connected components must all be of the same type. Thus, we consider two cases.
Case 1: $\mu_{2}\neq 0$ and each regular connected component is a deformed Sasakian structure. By Theorem D, each is a Riemannian submersion over a Kähler-Einstein manifold $(N,g_{N},J_{N})$. Since $N$ is of real dimension two, it must have constant curvature. Thus, the universal Sasakian model must be one of the three standard forms in dimension three [65]. The soliton equation (1.1) then reduces to ODEs which can be solved explicitly [30, 68] and, in particular, there is a singular orbit (two iff the manifold is compact).
Case 2: $\mu_{2}=0$ and each $\zeta_{c}$ is parallel with respect to the induced metric which is locally a Riemannian product. As in the proof of Theorem D, we write
$g$ $=dt^{2}+g_{t},$
$g_{t}$ $=H^{2}(t)\eta\otimes\eta+g^{\perp}_{t}.$
Since $\mu_{2}=0$, equations (2.5) and (4.1) implies that
$\frac{d}{dt}g^{\perp}_{t}=0.$
Thus, the metric is a local Riemannian product preserving the almost complex structure
$g=(dt^{2}+H^{2}(t)\eta\otimes\eta)+g^{\perp}.$
By the soliton system (2.7), $g^{\perp}$ has constant Ricci curvature and, thus, constant curvature (equal to $\lambda$) due to its low dimension. Also, the soliton equation restricted to first factor resembles one in dimension two: a warped product with warping function $|f^{\prime}|=|\nabla f|$. The analysis of that ODE is done in [6] and, when the curvature is non-constant, the completeness implies a singular critical level set consisting of a point at, without loss of generality, $t=0$. Accordingly, on $M$, the critical level set of $f$ at $t=0$ corresponds to a $2$-dimensional surface $N$ with the $t$-invariant metric $g^{\perp}$. Furthermore, each integral curve of $J(\nabla f)$ is homeomorphic to a circle. It follows that the soliton splits into a Riemannian product:
$(M_{1},dt^{2}+H^{2}(t)\eta\otimes\eta)\times(N,g_{N},J_{N}).$
When the curvature is constant, then $\mathrm{Hess}(f)$ is a constant multiple of the metric. Thus, for some constant $a$,
$f^{\prime}=\lambda t+a$
If $\lambda\neq 0$, then $f$ has a critical point and one argues as above to obtain a product structure. Otherwise, $\lambda=0$ and, consequently, the curvature of both $(dt^{2}+H^{2}(t)\eta\otimes\eta)$ and $g^{\perp}$ is vanishing. Then, $g$ is flat, a contradiction.
Proof of Corollary B.
By Theorem A, as the metric is irreducible, $(M,g,J,f)$ must be foliated by equidistant level sets of $f$ and a regular one is quotient of a connected deformed Sasakian structure and there is exactly one singular level set. By Theorem D, it is a Riemannian submersion, with circle fibers, over a Kähler-Einstein manifold $(N,g_{N},J_{N})$.
Claim: If $M$ is non-compact, then $N$ is simply connected.
Proof of the claim: Let $P$ denote the topological manifold of a regular level set. Since $M$ is non-compact, it is is constructed by collapsing at one end of $P\times[0,\infty]$. If the singular set is a point, the result follows immediately from Theorem D. If not, $M$ could be written as a union of an one-sided tubular neighborhood around the singular level set, diffeomorphic to $N$, and $P\times(0,\infty)$. The intersection is diffeomorphic to $P\times(0,\epsilon)$. It is immediate to construct a deformation retract from the tubular neighborhood to $N$ and from each of $P\times(0,\epsilon)$ and $P\times(0,\infty)$ to $P$. Consequently, Seifert-Van Kampen theorem deduces that, for $\pi_{1}$ denoting the fundamental group,
$\pi_{1}(M)=\pi_{1}(N)\ast_{\pi_{1}(P)}\pi_{1}(P).$
Since $M$ is simply connected, so is $N$.
Thus, $N$ is simply connected; $(N,g_{N},J_{N})$ is of constant curvature, and its isometry group is of dimension $2^{2}-1=3$. Using Theorem D and the simply-connectedness of $N$, the action lifts to one on each regular level set and combines with the circle action generated by $J(\nabla f)$ to forms an automorphism group on the deformed Sasakian structure; see [64, Lemma 5.1] and its refinement [12, Proposition 7.2.2 and Theorem 8.1.8]. Thus, each regular level set is a homogeneous deformed Sasakian structure and the isometry group of $(M,g,J,f)$ is precisely this automorphism group of dimension $3+1=4$.
Proof of Corollary C.
It follows from [56] that, for $\lambda\neq 0$ and irreducible GRS, the isometry group preserves the potential function $f$. The same conclusion also holds if the group is compact via an average argument as shown in [30].
As the structure is of cohomogeneity one, $f$ and $S\SS$ have same level sets and are functionally dependent. The result then follows from Theorem A. ∎
5.1. Statements and Declarations:
The work was partially supported by grants from the Simons Foundation [709791], the National Science Foundation [DMS-2104988], and the Vietnam Institute for Advanced Study in Mathematics.